Evaluate the function at the specified values of the indepent variable and simplify.

f(x)=x^4-4x^2+4
a) f(2)=
b) f(5)=
c) f(-1)=

Answer by fractalier(6550)

You merely plug the given values into the function
f(x)=x^4-4x^2+4
a) f(2)= 2^4 - 4(2^2) + 4 = 16 - 16 + 4 = 4
b) f(5)= 5^4 - 4(5^2) + 4 = 625 - 100 + 4 = 529
c) f(-1)= (-1)^4 - 4(-1)^2 + 4 = 1 - 4 + 4 = 1